OpenStudy (anonymous):

f(x) = { 3 x + 8 } /{ 5 x + 5 }, find f'( x ) find f'(3)

OpenStudy (austinl):

I assume you are referring to derivatives? \(\displaystyle \frac{d}{dx}(\frac{h(x)}{g(x)})=\frac{g(x)h^{\prime}(x)-h(x)g^{\prime}(x)}{g(x)^2}\) Does this make sense?

OpenStudy (austinl):

For your problem you have, \(\displaystyle f(x)=\frac{h(x)}{g(x)}=\frac{3x+8}{5x+5}\)

OpenStudy (anonymous):

Yep. I get the wrong answer when I try doing the derivitive

OpenStudy (austinl):

Ok, lets break this up shall we? \(\displaystyle h(x)=3x+8\) \(\displaystyle g(x)=5x+5\) \(h^{\prime}(x)=~?\) \(g^{\prime}(x)=~?\)

OpenStudy (austinl):

What is the derivative of 3x+8?

OpenStudy (anonymous):

3?

OpenStudy (austinl):

Yep, and how about 5x+5?

OpenStudy (anonymous):

5

OpenStudy (austinl):

Yeppers! So now we have, \(h(x)=3x+8\\ g(x)=5x+5\\ ~\\ h^{\prime}(x)=3\\ g^{\prime}(x)=5\) Now we can input those into this formula, \(\displaystyle \frac{d}{dx}(\frac{h(x)}{g(x)})=\frac{g(x)h^{\prime}(x)-h(x)g^{\prime}(x)}{g(x)^2}\) What would this look like?

OpenStudy (anonymous):

(5x+5)(3)-(3x+8)(5)/(5x+5)^2

OpenStudy (anonymous):

so then when you cancel out the (5x+5) from the numerator and denominator, it should be (3)-(3x+8) (5)/(5x+5) correct?

OpenStudy (anonymous):

Oh, i think i figured it out. The (5x+5) isn't allowed to cancel right?

OpenStudy (austinl):

\(\displaystyle \frac{(5x+5)(3)-(3x+8)(5)}{(5x+5)^2}\) This is a technically correct answer. I am sure you could multiply stuff out, and FOIL and such and get stuff to cancel.

OpenStudy (ranga):

{ (5x+5)(3)-(3x+8)(5) } / (5x+5)^2 Use distributive property and simplify the expression within the curly braces { } above.

OpenStudy (ranga):

@kquach1 : Your earlier cancelling out the (5x + 5) is not correct. You cancel it from the first expression which is correct. But the second expression does not have a (5x + 5) in the numerator and so the denominator will be (5x + 5)^2 for the second term. Instead just simplify the numerator in austinL's expression above and leave the denominator as (5x + 5)^2 or you can make the denominator 25(x + 1)^2. That will be your f'(x). Substitute x = 3 to evaluate f'(3).